OpenStudy (anonymous):

how many solutions do you get from this system of equations. y=3(x-1) -y+3x=3

OpenStudy (mathstudent55):

Can you solve both equations for y and simplify the right side?

OpenStudy (anonymous):

me

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you please help i dont understand it AT ALL. Please help i will give medal. @mathstudent55

OpenStudy (mathstudent55):

Let's take the first equation: y = 3(x - 1) We can simplify the right side using the distributive property. y = 3x - 3 We stop here with the first equation. The second equation is: -y + 3x = 3

OpenStudy (mathstudent55):

Now we solve the second equation for y. We subtract 3x from both sides: -y = -3x + 3 Now we multiply both sides by -1: y = 3x - 3 We stop here with the second equation.

OpenStudy (mathstudent55):

Now we look at both equations: y = -3x + 3 y = -3x + 3 Both equations are the same equation. When this happens, there is an infinite number of solutions. Any point on line y = -3x + 3 is a solution, and there are infinite points on the line.